OpenStudy (anonymous):

Null factor of: 4x^2 = 2x

OpenStudy (anonymous):

From this we rearrange: \[0=4x^2-2x\] \[0=2x(2x-1)\] Therefore: 2x = 0, or 2x-1 = 0 Giving: x = 0, or x = 1/2 Is that what you were looking for?

OpenStudy (anonymous):

thank you